a distance of about 50 cm is measured using a metre stick having mm divisison.the %error is??how is %error calculated??

Grade:

1 Answers

Ramesh V
14 years ago

if the least count of scale is 1 mm

and let say the scale measured the distance as = 50.5cms

original distance is 50 cms

so % error is : (50.5 - 50) /50 =  + 1 %

if it measures say 49.5cms, then error is = (49.5 - 50)/50 = -1%

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult
to understand - post it here and we will get you the answer and detailed solution very quickly.We are all
IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Ramesh
IIT Kgp - 05 batch